# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ Acknowledgements
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 850930). The second author was supported by the LABEX MILYON of Université de Lyon. The authors would like to thank Damien Gaboriau, Dominik Kirstein, and Ric Wade for helpful conversations.
@@ 2. Background on the Farrell–Jones conjecture
For full context and background on the Farrell–Jones conjecture, see for instance Lück’s book project [Lüc]. In this section, we recall some properties of the class $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ of groups which satisfy the Farrell–Jones conjecture for $X$-theory where $X$ is $A$, $K$, or $L$.
@@
The cases of $K$- and $L$-theory are given in [GMR15, Theorem 2.1]. The case of $A$-theory is [${}^{+}$18, Theorem 1.1(ii)]. ∎
@@
The cases of $K$- and $L$-theory are given in [BFL14, Theorem 1.7]. The case of $A$-theory is given in [${}^{+}$18, Theorem 1.1(ii)]. ∎
@@
For the first statement apply 2.2 to the short exact sequence, noting that both finite groups and virtually cyclic groups are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. For the second, observe that if $\Gamma$ is residually finite then there is a finite index subgroup $\Gamma_{0}$ of $\Gamma$ whose intersection with $N$ is trivial, and then $\Gamma_{0}\cong Q_{0}$ for some finite index subgroup $Q_{0}$ of $Q$. The result follows from commensurability. ∎
We refer for instance to the work of Bowditch [Bow12] for the definition of a relatively hyperbolic group.
@@
This result is due to Bartels. The cases of $K$- and $L$-theory are [Bar17, Corollary 4.6]. The case of $A$-theory is also ostensibly due to Bartels combined with some recent developments on the $A$-theoretic FJC. We sketch the relevant details. The key here is that Bartels’ space $\Delta$ for a relatively hyperbolic group pair $(G,{\mathcal{P}})$ is finitely ${\mathcal{P}}$-amenable (see [Bar17, Theorem 3.1]). By [Kno19, Proof of Theorem 1.8(a)] this implies that $G$ is strongly transfer reducible over ${\mathcal{F}}$. The result now follows from [${}^{+}$18, Theorem 6.19]. ∎
Let $G$ be a group acting by isometries on a tree $T$. Recall that the action is acylindrical if there exists $K\geq 0$ such that the stabiliser of any geodesic path of length at least $K$ is finite.
@@
The result is due to S. Knopf. For $K$-theory we refer to [Kno19, Corollary 4.2]. The result for $L$-theory is [Kno19, Corollary 4.3], note that here one has the additional hypothesis that index $2$ overgroups of the stabilisers in $G$ must satisfy $\operatorname{\mathbf{FJC}}_{\mathbf{L}}$. But this follows from 2.1. For $A$-theory, as in Bartels’ result, one combines finite ${\mathcal{F}}$-amenability [Kno19, Proposition 4.1] with the recent developments for $A$-theory [Kno19, Proof of Theorem 1.8(a)] and [${}^{+}$18, Theorem 6.19]. ∎
Let $G$ be a group and let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. Let $\mathrm{Per}(\Phi)=\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\in\mathbb{N}}$ be the periodic subgroup of $\Phi$. At several points in our arguments we will need the following lemma.
@@
$\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}=\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\in\mathbb{N}}\rtimes_{\Phi}\mathbb{Z}=\bigcup_{n\in\mathbb{N}}\left(\operatorname{\mathrm{Fix}}(\Phi^{n!})\rtimes_{\Phi}\mathbb{Z}\right),$
@@ 3.A. Free products of groups
Let $N\in\mathbb{N}$, let $G_{1},\ldots,G_{k}$ be countable groups and let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$. Let =${\mathcal{F}}=\{[G_{1}],\ldots,[G_{k}]\}$ be the set consisting of the conjugacy classes of the $G_{i}$. We refer to $(G,{\mathcal{F}})$ as a free product.
An element $g\in G$ is peripheral if there exists $[A]\in{\mathcal{F}}$ with $g\in A$. Otherwise, $g$ is nonperipheral. A subgroup $P$ of $G$ is peripheral if every element of $P$ is peripheral, and is nonperipheral otherwise.
A free factor system of $(G,{\mathcal{F}})$ is a set =${\mathcal{F}}^{\prime}=\{[A_{1}],\ldots,[A_{\ell}]\}$ of conjugacy classes of proper subgroups of $G$ such that:
@@
    for every $i\in\{1,\ldots,k\}$, there exists $[A]\in{\mathcal{F}}^{\prime}$ such that $G_{i}\subseteq A$;
@@
The set of free factor systems of $G$ is equipped with a partial order where ${\mathcal{F}}_{1}\leq{\mathcal{F}}_{2}$ if, for every $[A_{1}]\in{\mathcal{F}}_{1}$, there exists $[A_{2}]\in{\mathcal{F}}_{2}$ with $A_{1}\subseteq A_{2}$. A free factor system ${\mathcal{F}}^{\prime}$ is sporadic if either =${\mathcal{F}}^{\prime}=\{[A_{1}],[A_{2}]\}$ and $G=A_{1}\ast A_{2}$ or =${\mathcal{F}}^{\prime}=\{[A_{1}]\}$ and $G=A_{1}\ast\mathbb{Z}$. Otherwise, the free factor system ${\mathcal{F}}^{\prime}$ is nonsporadic. The free product $(G,{\mathcal{F}})$ is sporadic (resp. nonsporadic) if ${\mathcal{F}}$ is.
We denote by $\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ the subgroup of automorphisms of $G$ preserving ${\mathcal{F}}$ and by $\operatorname{\mathrm{Out}}(G,{\mathcal{F}})$ the subgroup of outer automorphisms of $G$ preserving ${\mathcal{F}}$. An automorphism $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ is fully irreducible if no power of $\Phi$ fixes a free factor system of $(G,{\mathcal{F}})$.
A $(G,{\mathcal{F}})$-tree is a tree equipped with an action of $G$ without inversion such that, for every $[A]\in{\mathcal{F}}$, the group $A$ is elliptic in $T$. A Grushko $(G,{\mathcal{F}})$-tree is a $(G,{\mathcal{F}})$-tree $T$ with trivial edge stabilisers and such that, for every $v\in VT$, the conjugacy class of the stabiliser $G_{v}$ of $v$ is trivial or contained in ${\mathcal{F}}$.
Let ${\mathcal{F}}^{\prime}$ be a sporadic free factor system of $(G,{\mathcal{F}})$. There is a unique, up to unique $G$-equivariant homeomorphism, reduced $(G,{\mathcal{F}}^{\prime})$-tree $T_{{\mathcal{F}}^{\prime}}$, which we call the Bass-Serre tree of $(G,{\mathcal{F}}^{\prime})$. The tree $T_{{\mathcal{F}}^{\prime}}$ has a unique orbit of edges. The tree $T_{{\mathcal{F}}^{\prime}}$ is canonical in the sense that every element $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$ induces a $G$-equivariant homeomorphism of $T_{{\mathcal{F}}^{\prime}}$. Therefore, for every $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$, the group $G\rtimes_{\Phi}\mathbb{Z}$ acts by homeomorphisms on $T_{{\mathcal{F}}^{\prime}}$.
@@ 3.B. Growth under an automorphism of a free product
Let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. An element $g\in G$ has .$\lVert.\rVert_{T}$-polynomial growth under iteration of $\Phi$ if there exists $P\in\mathbb{Z}[X]$ such that, for every $n\in\mathbb{N}$:
@@
Note that any elliptic element of $G$ in $T$ has .$\lVert.\rVert_{T}$-polynomial growth under iteration of $\Phi$.
A subgroup $P$ of $G$ is a .$\lVert.\rVert_{T}$-polynomial subgroup of $\Phi$ if there exists an automorphism $\Psi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ contained in the outer class of some power of $\Phi$ such that $\Psi(P)=P$ and every element of $P$ has .$\lVert.\rVert_{T}$-polynomial growth under iteration of $\Psi$.
Let ${\mathcal{P}}_{T}(\Phi)$ be the set of conjugacy classes of maximal .$\lVert.\rVert_{T}$-polynomial subgroups of $\Phi$. When $\Phi$ is fully irreducible, the set ${\mathcal{P}}_{T}(\Phi)$ satisfies some additional properties. Recall that a subgroup $A$ of $G$ is malnormal if, for every $g\in G-A$, we have $A\cap gAg^{-1}=\{e\}$.
@@
[DL22, Proposition 1.13] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. Let $T$ be a Grushko $(G,{\mathcal{F}})$-tree.
@@
    For every $[A]\in{\mathcal{P}}_{T}(\Phi)$, the subgroup $A$ is malnormal in $G$.
@@
The following result is due to Dahmani–Li [DL22].
@@
[DL22, Corollary 2.3] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. Let $T$ be a Grushko $(G,{\mathcal{F}})$-tree and let ${\mathcal{P}}_{T}(\Phi)$ be the set of conjugacy classes of maximal .$\lVert.\rVert_{T}$-polynomial subgroups of $\Phi$. There exists $n\in\mathbb{N}$ such that the group $G\rtimes_{\Phi^{n}}\mathbb{Z}$ is hyperbolic relative to the suspension of ${\mathcal{P}}_{T}(\Phi)$.
Note that Theorem 3.2 implies that the set ${\mathcal{P}}_{T}(\Phi)$ does not depend on $T$ when $\Phi$ is fully irreducible. In the rest of the section, we give a precise description of the set ${\mathcal{P}}_{T}(\Phi)$ for a fully irreducible automorphism.
We first need a result, which can be found for instance in the work of Francaviglia–Martino–Syrigos [FMS21] concerning the existence of a limiting tree of a fully irreducible automorphism.
@@
[FMS21, Lemma 2.14.1] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be a fully irreducible automorphism. There exist a Grushko $(G,{\mathcal{F}})$-tree $S$, an $\mathbb{R}$-tree $T$ equipped with an isometric action of $G$ and a constant $\lambda>1$ such that, for every $g\in G$, we have
@@
Let $(G,{\mathcal{F}})$ be a free product. Note that, for any subgroup $A$ of $G$, the free factor system ${\mathcal{F}}$ induces a free factor system ${\mathcal{F}}|_{A}$ of $A$. Following the terminology of for instance Guirardel-Horbez [GH22, Definition 3.2], an $\mathbb{R}$-tree equipped with an isometric action of $G$ is an arational $(G,{\mathcal{F}})$-tree if the following holds:
@@
    for every $[A]\in{\mathcal{F}}$, the group $A$ is elliptic in $T$;
@@
    for every free factor system ${\mathcal{F}}<{\mathcal{F}}^{\prime}$ and every $[A]\in{\mathcal{F}}^{\prime}$ such that $A$ is nonperipheral, the action of $A$ on its minimal tree in $T$ is a Grushko $(A,{\mathcal{F}}|_{A})$-tree.
@@
Let $(G,{\mathcal{F}})$ be a nonsporadic free product, let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be a fully irreducible automorphism and let $S$ be a Grushko $(G,{\mathcal{F}})$-tree given by Lemma 3.3. For every $[P]\in{\mathcal{P}}_{S}(\Phi)$, either $[P]\in{\mathcal{F}}$ or $P$ is nonperipheral and infinite cyclic.
@@
Let $T$ be the $\mathbb{R}$-tree associated with $S$ given by Lemma 3.3. Note that, by Equation (1), for every .$\lVert.\rVert_{S}$-polynomially growing element $g\in G$, we have $\lVert g\rVert_{T}=0$.
By [GH22, Theorems 3.4,4.1] the $\mathbb{R}$-tree $T$ is an arational $(G,{\mathcal{F}})$-tree. It has trivial arc stabilisers (because it is mixing [Hor14, Lemma 4.9]). Thus, for every $[P]\in{\mathcal{P}}_{S}(\Phi)$, the group $P$ fixes a point in $T$.
By [Hor14, Lemma 4.6], using the fact that $T$ is arational, for every point $x\in T$, the stabiliser $G_{x}$ of $x$ is either peripheral or nonperipheral and infinite cyclic. Thus, for every $[P]\in{\mathcal{P}}_{S}(\Phi)$, the elliptic subgroup $P$ is either peripheral or nonperipheral and infinite cyclic. By maximality of $P$, either $[P]\in{\mathcal{F}}$ or $P$ is nonperipheral and infinite cyclic. ∎

### 02 Mar 2026 v2

@@ Acknowledgements
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 850930). The second author was supported by the LABEX MILYON of Université de Lyon. The authors would like to thank Damien Gaboriau, Dominik Kirstein, Gilbert Levitt, and Ric Wade for helpful conversations. The authors would also like to thank the anonymous referee for a number of helpful comments and clarifications and for suggesting an elegant simplification of some of our arguments.
@@ 2. Background on the Farrell–Jones conjecture
For full context and background on the Farrell–Jones conjecture, see for instance Lück’s book project [41]. In this section, we recall some properties of the class $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ of groups which satisfy the Farrell–Jones conjecture for $X$-theory where $X$ is $A$, $K$, or $L$.
@@
The cases of $K$- and $L$-theory are given in [22, Theorem 2.1]. The case of $A$-theory is [18, Theorem 1.1(ii)]. ∎
@@
The cases of $K$- and $L$-theory are given in [1, Theorem 1.7]. The case of $A$-theory is given in [18, Theorem 1.1(ii)]. ∎
@@
For the first statement apply ˜2.2 to the short exact sequence, noting that both finite groups and virtually cyclic groups are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. For the second, observe that if $\Gamma$ is residually finite then there is a finite index subgroup $\Gamma_{0}$ of $\Gamma$ whose intersection with $N$ is trivial, and then $\Gamma_{0}\cong Q_{0}$ for some finite index subgroup $Q_{0}$ of $Q$. The result follows from commensurability. ∎
We refer for instance to the work of Bowditch [13] for the definition of a relatively hyperbolic group.
@@
This result is due to Bartels. The cases of $K$- and $L$-theory are [8, Corollary 4.6]. The case of $A$-theory is also ostensibly due to Bartels combined with some recent developments on the $A$-theoretic FJC. We sketch the relevant details. The key here is that Bartels’ space $\Delta$ for a relatively hyperbolic group pair $(G,{\mathcal{P}})$ is finitely $\mathcal{P}$}-amenable (see [8, Theorem 3.1]). By [37, Proof of Theorem 1.8(a)] this implies that $G$ is strongly transfer reducible over $\mathcal{F}$}. The result now follows from [18, Theorem 6.19]. ∎
Let $G$ be a group acting by isometries on a tree $T$. Recall that the action is acylindrical if there exists $K\geq 0$ such that the stabiliser of any geodesic path of length at least $K$ is of bounded size.
@@
The result is due to S. Knopf. For $K$-theory we refer to [37, Corollary 4.2]. The result for $L$-theory is [37, Corollary 4.3], note that here one has the additional hypothesis that index $2$ overgroups of the stabilisers in $G$ must satisfy $\operatorname{\mathbf{FJC}}_{\mathbf{L}}$. But this follows from ˜2.1. For $A$-theory, as in Bartels’ result, one combines finite $\mathcal{F}$}-amenability [37, Proposition 4.1] with the recent developments for $A$-theory [37, Proof of Theorem 1.8(a)] and [18, Theorem 6.19]. ∎
Let $G$ be a group and let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. Let $\mathrm{Per}(\Phi)=\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\geq 1}$ be the periodic subgroup of $\Phi$. At several points in our arguments we will need the following lemma.
@@
$\mathrm{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}=\langle\operatorname{\mathrm{Fix}}(\Phi^{n})\rangle_{n\geq 1}\rtimes_{\Phi}\mathbb{Z}=\bigcup_{n\geq 1}\left(\operatorname{\mathrm{Fix}}(\Phi^{n!})\rtimes_{\Phi}\mathbb{Z}\right),$
@@ 3.A. Free products of groups
Let $N\in\mathbb{N}$, let $G_{1},\ldots,G_{k}$ be countable groups and let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$. Let ${\mathcal{F}}=\{[G_{1}],\ldots,[G_{k}]\}$ be the set consisting of the conjugacy classes of the $G_{i}$. We refer to $(G,{\mathcal{F}})$ as a free product.
An element $g\in G$ is peripheral if there exists $A]\in{\mathcal{F}}$ with $g\in A$. Otherwise, $g$ is nonperipheral. A subgroup $P$ of $G$ is peripheral if every element of $P$ is peripheral, and is nonperipheral otherwise.
A free factor system of $(G,{\mathcal{F}})$ is a set ${\mathcal{F}}^{\prime}=\{[A_{1}],\ldots,[A_{\ell}]\}$ of conjugacy classes of proper subgroups of $G$ such that:
@@
    for every $i\in\{1,\ldots,k\}$, there exists $A]\in{\mathcal{F}}^{\prime}$ such that $G_{i}\subseteq A$;
@@
The set of free factor systems of $G$ is equipped with a partial order where ${\mathcal{F}}_{1}\leq{\mathcal{F}}_{2}$ if, for every $A_{1}]\in{\mathcal{F}}_{1}$, there exists $A_{2}]\in{\mathcal{F}}_{2}$ with $A_{1}\subseteq A_{2}$. A free factor system $\mathcal{F}$$^{\prime}$ is sporadic if either ${\mathcal{F}}^{\prime}=\{[A_{1}],[A_{2}]\}$ and $G=A_{1}\ast A_{2}$ or ${\mathcal{F}}^{\prime}=\{[A_{1}]\}$ and $G=A_{1}\ast\mathbb{Z}$. Otherwise, the free factor system $\mathcal{F}$$^{\prime}$ is nonsporadic. The free product $(G,{\mathcal{F}})$ is sporadic (resp. nonsporadic) if $\mathcal{F}$} is.
We denote by $\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ the subgroup of automorphisms of $G$ preserving $\mathcal{F}$} and by $\operatorname{\mathrm{Out}}(G,{\mathcal{F}})$ the subgroup of outer automorphisms of $G$ preserving $\mathcal{F}$}. An automorphism $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ is fully irreducible if no power of $\Phi$ fixes a proper free factor system of $(G,{\mathcal{F}})$.
A $(G,{\mathcal{F}})$-tree is a tree equipped with an action of $G$ without inversion such that, for every $A]\in{\mathcal{F}}$, the group $A$ is elliptic in $T$. A Grushko $(G,{\mathcal{F}})$-tree is a $(G,{\mathcal{F}})$-tree $T$ with trivial edge stabilisers and such that, for every $v\in VT$, the conjugacy class of the stabiliser $G_{v}$ of $v$ is trivial or contained in $\mathcal{F}$}.
Let $\mathcal{F}$$^{\prime}$ be a sporadic free factor system of $(G,{\mathcal{F}})$. There is a unique, up to unique $G$-equivariant homeomorphism, reduced Grushko $(G,{\mathcal{F}}^{\prime})$-tree $T_{{\mathcal{F}}^{\prime}}$, which we call the Bass-Serre tree of $(G,{\mathcal{F}}^{\prime})$. (In this context, reduced means that no vertices have trivial stabiliser.) The tree $T_{{\mathcal{F}}^{\prime}}$ has a unique orbit of edges. The tree $T_{{\mathcal{F}}^{\prime}}$ is canonical in the sense that every element $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$ induces a $G$-equivariant homeomorphism of $T_{{\mathcal{F}}^{\prime}}$. Therefore, for every $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$, the group $G\rtimes_{\Phi}\mathbb{Z}$ acts by homeomorphisms on $T_{{\mathcal{F}}^{\prime}}$.
@@ 3.B. Growth under an automorphism of a free product
Let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. An element $g\in G$ has $\lVert.\rVert_{T}$-polynomial growth under iteration of $\Phi$ if there exists $P\in\mathbb{Z}[X]$ such that, for every $n\in\mathbb{N}$:
@@
Note that any elliptic element of $G$ in $T$ has $\lVert.\rVert_{T}$-polynomial growth under iteration of $\Phi$.
A subgroup $P$ of $G$ is a $\lVert.\rVert_{T}$-polynomial subgroup of $\Phi$ if there exists an automorphism $\Psi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ contained in the outer class of some power of $\Phi$ such that $\Psi(P)=P$ and every element of $P$ has $\lVert.\rVert_{T}$-polynomial growth under iteration of $\Psi$.
Let ${\mathcal{P}}_{T}(\Phi)$ be the set of conjugacy classes of maximal $\lVert.\rVert_{T}$-polynomial subgroups of $\Phi$. By for instance [20, Theorem 1.1], the set ${\mathcal{P}}_{T}(\Phi)$ does not depend on $T$.
When $\Phi$ is fully irreducible, the set ${\mathcal{P}}_{T}(\Phi)$ satisfies some additional properties. Recall that a subgroup $A$ of $G$ is malnormal if, for every $g\in G-A$, we have $A\cap gAg^{-1}=\{e\}$.
@@
[16, Proposition 1.13] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. Let $T$ be a Grushko $(G,{\mathcal{F}})$-tree.
@@
    For every $A]\in{\mathcal{P}}_{T}(\Phi)$, the subgroup $A$ is malnormal in $G$.
@@
The following result is due to Dahmani–Li [16].
@@
[16, Corollary 2.3] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be fully irreducible. Let $T$ be a Grushko $(G,{\mathcal{F}})$-tree and let ${\mathcal{P}}_{T}(\Phi)$ be the set of conjugacy classes of maximal $\lVert.\rVert_{T}$-polynomial subgroups of $\Phi$. There exists $n\in\mathbb{N}$ such that the group $G\rtimes_{\Phi^{n}}\mathbb{Z}$ is hyperbolic relative to the suspension of ${\mathcal{P}}_{T}(\Phi)$.
In the rest of the section, we give a precise description of the set ${\mathcal{P}}_{T}(\Phi)$ for a fully irreducible automorphism.
We first need a result, which can be found for instance in the work of Francaviglia–Martino–Syrigos [21] concerning the existence of a limiting tree of a fully irreducible automorphism.
@@
[21, Lemma 2.14.1] Let $(G,{\mathcal{F}})$ be a nonsporadic free product and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be a fully irreducible automorphism. There exist a Grushko $(G,{\mathcal{F}})$-tree $S$, an $\mathbb{R}$-tree $T$ equipped with an isometric action of $G$ and a constant $\lambda>1$ such that, for every $g\in G$, we have
@@
Let $(G,{\mathcal{F}})$ be a free product. Note that, for any subgroup $A$ of $G$, the free factor system $\mathcal{F}$} induces a free factor system ${\mathcal{F}}|_{A}$ of $A$. Following the terminology of for instance Guirardel-Horbez [24, Definition 3.2], an $\mathbb{R}$-tree equipped with an isometric action of $G$ is an arational $(G,{\mathcal{F}})$-tree if the following holds:
@@
    for every $A]\in{\mathcal{F}}$, the group $A$ is elliptic in $T$;
@@
    for every free factor system $\mathcal{F}$$\mathcal{F}$$^{\prime}$ and every $A]\in{\mathcal{F}}^{\prime}$ such that $A$ is nonperipheral, the action of $A$ on its minimal tree in $T$ is a Grushko $(A,{\mathcal{F}}|_{A})$-tree.
@@
Let $(G,{\mathcal{F}})$ be a nonsporadic free product, let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$ be a fully irreducible automorphism and let $S$ be a Grushko $(G,{\mathcal{F}})$-tree given by Lemma 3.3. For every $P]\in{\mathcal{P}}_{S}(\Phi)$, either $P]\in{\mathcal{F}}$ or $P$ is nonperipheral and infinite cyclic.
@@
Let $T$ be the $\mathbb{R}$-tree associated with $S$ given by Lemma 3.3. Note that, by Equation (1), for every $\lVert.\rVert_{S}$-polynomially growing element $g\in G$, we have $\lVert g\rVert_{T}=0$.
By [24, Theorems 3.4,4.1] the $\mathbb{R}$-tree $T$ is an arational $(G,{\mathcal{F}})$-tree. It has trivial arc stabilisers (because it is mixing [32, Lemma 4.9]). Thus, for every $P]\in{\mathcal{P}}_{S}(\Phi)$, the group $P$ fixes a point in $T$.
By [32, Lemma 4.6], using the fact that $T$ is arational, for every point $x\in T$, the stabiliser $G_{x}$ of $x$ is either peripheral or nonperipheral and infinite cyclic. Thus, for every $P]\in{\mathcal{P}}_{S}(\Phi)$, the elliptic subgroup $P$ is either peripheral or nonperipheral and infinite cyclic. By maximality of $P$, either $P]\in{\mathcal{F}}$ or $P$ is nonperipheral and infinite cyclic. ∎